OpenStudy (anonymous):

Cynthia has a cup full of dimes and quarters that she would like to put in her savings account. When she takes the cup to the bank, the banker tells Cynthia that the 386 coins in the cup are worth $58.10. How many quarters does Cynthia have in her cup? A. 126 B. 130 C. 256 D. 386

OpenStudy (anonymous):

Solve 386=a+b a is quarters b is dimes Then do 58.16=.25a+.1b solve for b or a in first equation then plug into second equation to get ur answer.

OpenStudy (anonymous):

What do I plugin?

OpenStudy (anonymous):

Set actually b = 386 - a because u need to set the second equation with only one unknown and that's quarters. So then u have .25a-.1(386-a)= 58.16

OpenStudy (anonymous):

Would I get A?

OpenStudy (anonymous):

You are solving two equations for two unknowns u can't have two unknowns in one equation. Solve ur second equation for (a) because u need to find quarters.

OpenStudy (anonymous):

But when u solve the second equation u solve it only for variable (a). Ur variable b is solved from the first equation so you are left with 130.

OpenStudy (anonymous):

Okay, so I got B.

OpenStudy (anonymous):

Is that right?

OpenStudy (anonymous):

I messed up my steps hold on you want to solve your second equation in terms of dimes u should get a amount of coins. Then to solve for your quarters you plug ur coins into the first equation so then it is 386 - dimes = quarters.

OpenStudy (anonymous):

Ugh, you're confusing me.

OpenStudy (anonymous):

The letter a is the amount of quarters. Solve the equation 58.16=.25a+ .1(386-a) when you solve it for a you will get 130. Which is the amount of quarters you have.

OpenStudy (anonymous):

Alright, Thanks! Do you mind helping with this one as well? A local florist charges $8 for a red rose and $10 for a white rose. He sold twice as many red roses as white roses for a total income of $8,554 for the day. How many roses did he sell all together? A. 329 B. 658 C. 987 D. 8,554

OpenStudy (anonymous):

he sold 987 roses